$SPAD/src/input frac.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool frac.output
)set message test on
)set message auto off
)clear all
--S 1 of 12
a := 11/12
--R
--R
--R        11
--R   (1)  --
--R        12
--R                                                      Type: Fraction(Integer)
--E 1
--S 2 of 12
b := 23/24
--R
--R
--R        23
--R   (2)  --
--R        24
--R                                                      Type: Fraction(Integer)
--E 2
--S 3 of 12
3 - a*b**2 + a + b/a
--R
--R
--R        313271
--R   (3)  ------
--R         76032
--R                                                      Type: Fraction(Integer)
--E 3
--S 4 of 12
numer(a)
--R
--R

--R   (4)  11
--R                                                        Type: PositiveInteger
--E 4
--S 5 of 12
denom(b)
--R
--R
--R   (5)  24
--R                                                        Type: PositiveInteger
--E 5
--S 6 of 12
r := (x**2 + 2*x + 1)/(x**2 - 2*x + 1)
--R
--R
--R         2
--R        x  + 2x + 1
--R   (6)  -----------
--R         2
--R        x  - 2x + 1
--R                                          Type: Fraction(Polynomial(Integer))
--E 6
--S 7 of 12
factor(r)
--R
--R
--R         2
--R        x  + 2x + 1
--R   (7)  -----------
--R         2
--R        x  - 2x + 1
--R                                Type: Factored(Fraction(Polynomial(Integer)))
--E 7
--S 8 of 12
map(factor,r)
--R
--R
--R               2
--R        (x + 1)
--R   (8)  --------
--R               2
--R        (x - 1)
--R                                Type: Fraction(Factored(Polynomial(Integer)))
--E 8
--S 9 of 12
continuedFraction(7/12)

--R
--R
--R          1 |     1 |     1 |     1 |
--R   (9)  +---+ + +---+ + +---+ + +---+
--R        | 1     | 1     | 2     | 2
--R                                             Type: ContinuedFraction(Integer)
--E 9
--S 10 of 12
partialFraction(7,12)
--R
--R
--R              3   1
--R   (10)  1 - -- + -
--R              2   3
--R             2
--R                                               Type: PartialFraction(Integer)
--E 10
--S 11 of 12
g := 2/3 + 4/5*%i
--R
--R
--R         2   4
--R   (11)  - + - %i
--R         3   5
--R                                             Type: Complex(Fraction(Integer))
--E 11
--S 12 of 12
g :: FRAC COMPLEX INT
--R
--R
--R         10 + 12%i
--R   (12)  ---------
--R             15
--R                                             Type: Fraction(Complex(Integer))
--E 12
)spool
)lisp (bye)

———-

References^{[1] nothing}

6